Activity 1:


Solve the following problems applying the concepts learned above. Write your answer on a separate sheet of paper. (Show your solution)


1. The rate of change x is proportional to x. When t =0, x0 = 3 and when t =2, x=6. What is the value of x when t =4?


2. A certain plutonium isotope decays at a rate proportional to the amount present. Approximately 15% of the original amount decomposes in 100 years. How much amount of the substance has decayed after 600 years? Also, find the half - life t1/2 of this radioactive substance ; that is, find the time required for this substance to decay to one- half of its original amount.




1
Expert's answer
2022-05-09T16:08:29-0400

1.


"dx\/dt=kx"

"\dfrac{dx}{x}=kdt"

"\int \dfrac{dx}{x}=\int kdt"

"\ln |x|=kt+\ln C"

"x=Ce^{kt}"

Given "x(0)=3"


"3=Ce^{0}=>C=3"

"x=3e^{kt}"

Given "x(2)=6"


"6=3e^{k(2)}"

"2k=\ln 2"

"k=\dfrac{\ln 2}{2}"

"x=3(2)^{t\/2}"

Then


"x(4)=3(2)^{4\/2}"

"x(4)=12"

2.


"N=\dfrac{1}{k}\dfrac{dN}{dt}"

"\dfrac{dN}{N}=kdt"

"\int \dfrac{dN}{N}=\int kdt"

"\ln |N|=kN+\ln C"

"N(t)=N_0e^{kt}"

Given "N(100)=0.85N_0"


"0.85N_0=N_0e^{100k}"

"k=0.01\ln0.85"

"N(t)=N_0(0.85)^{t\/100}"

Then


"N(600)=N_0(0.85)^{600\/100}"

"N(600)=0.37715N_0"

"100-37.715=62.285"

Approximately 62.285% of the original amount decomposes in 600 years.



"N(t_{1\/2})=N_0\/2=N_0(0.85)^{t_{1\/2}\/100}"

"t_{1/2}=100(\dfrac{\ln0.5}{\ln0.85})"

"t_{1/2}=426.5\ years"